### v2

@@ 4.2.4 Task Encoder
Figure 10(d) illustrates the task encoder, designed to extract task-specific embedding features from the multimodal token embeddings, given the special $<TASK>$ token embedding. The task encoder was implemented with multiple Transformer layers and regards all tokens as a single input sequence. Note that only the special $<TASK>$ token is forwarded to the task encoder and the rest of the question tokens are ignored, as we empirically found that other tokens are not useful in practice. The special token embedding from the final Transformer layer serves as the task-specific embedding feature that integrates all multimodal data. For M3FM-Base, M3FM-Large, and M3FM-Huge, the number of Transformer layers is 4 in every case, and the sizes of task token embeddings are 768, 1,024, and 1,280, respectively.
@@ 4.3 Predictors
The predictors map task-specific embedding features to answers. We implemented all predictors as a two-layer MLP. Different tasks may have different predictors or shared predictors for the same output dimension, such as ”Yes” or ”No” answers. Similar to language models that regard text generation as the token classification over a vocabulary problem, we formulate our answer prediction as a classification problem over the predefined answer candidates, as summarized in Table 1, except for six-year lung cancer risk prediction. Following [12], we formulate lung cancer risk prediction as a hazard regression problem. It is worth mentioning that more types of prediction tasks, such as image segmentation and object detection, can be incorporated into M3FM by adding the corresponding light-weight task-specific predictors as demonstrated in our previous study [58].
@@ 4.4 Self-supervised Pretraining
A key step to optimize large models is self-supervised pretraining with large unlabeled datasets. In this study, we adapted the masked autoencoder method [53, 54] to pretrain our CTViT on the constructed OpenM3Chest pretraining dataset. The image encoder introduced in Subsection 4.2.2 was optimized by predicting masked cubes (85%) from a small number of visible cubes (15%) using an additional shallow transformer decoder. To reduce the memory overhead, only some selected slices along the longitude direction were predicted while recovering each 3D patch. We pretrained CTViT with the defined multi-scale 3D CT volumes and a set of data augmentation operations, including random cropping, rotation, resizing, and perturbed display windowing. The text Transformer in M3FM was initialized with the off-the-shelf RoBERTa model pretrained via masked language modeling and then trained end-to-end [59, 52].
@@ 4.5 Multitask Learning
After self-supervised pretraining, M3FM can be trained with any combination of different tasks with properly selected multimodal datasets by optimizing multitask loss functions simultaneously. We use the softmax cross-entropy loss function for the classification task and the sigmoid cross-entropy loss function for the regression task. As the number of tasks increases, there is a significant rise in computational cost. To address this problem, we design a distributed task-parallel (DTP) training strategy. TDP assigns each computing device with a single task and a single data loader while the total number of training samples remains fixed across all devices for each task. Since M3FM is a unified model capable of handling various tasks, despite differences in input and output dimensions, gradients computed across all tasks can be readily accumulated, enabling simultaneous parameter optimization.
@@ 4.6 Transfer Learning
M3FM is designed for adaptability and generalization, enabling the enhancement of out-of-distribution task performance through transfer learning. This capability extends to new tasks with varying image input dimensions, clinical data types, and output dimensions. To accommodate different image dimensions, the addition of a linear embedding layer suffices. For diverse clinical datasets, we can simply describe involved clinical data in free text to the model, without needing any modification on the M3FM architecture. Specifically, adjusting to different output dimensions requires only the inclusion of a lightweight predictor. Consequently, M3FM can be easily fine-tuned to enable new tasks by leveraging the pre-trained model parameters.
4.7 Summary of Novel Methodological Elements
Multiple novel methodological elements are synergistically integrated to enable the excellent performance of our M3FM. The following key points are instrumental.
(1) The whole systematic workflow has been streamlined for the development of a hospital-grade foundation model, from clinical multitask definition to multimodal data curation, from radiologists’ reading procedure to the unified M3FM framework, from self-supervised pre-training to multitask learning with high-dimensional multimodal datasets. Importantly, the whole pipeline is designed with scalability in mind, allowing M3FM to be readily scaled up by integrating more training datasets and undertaking a broader range of clinical tasks.
(2) M3FM is the first medical multimodal multitask foundation model that effectively encodes different combinations of multimodal datasets, especially high-dimensional tomographic images at different scales—, and flexibly performs multiple tasks via free-text prompting. M3FM sets new performance records for a variety of tasks in the LDCT LCS process and beyond.
(3) The CT Vision Transformer (CTViT) has been invented that is compatible with multiple image sizes as an image encoder of M3FM. Our dedicated multi-scale CT tokenizer adaptively decomposes different sizes of volumetric CT inputs to facilitate task-specific analysis. CTViT disentangles the physical size from CT images and embeds it as a separate input. This allows M3FM to accommodate different scales of CT images without losing their physical size information, which helps specific tasks significantly, as demonstrated in Sub-section 2.3.
(4) The multi-scale volumetric CT self-supervised pretraining method has been proposed to pretrain M3FM, aided by the masked autoencoder.
(5) The distributed task-parallel (DTP) training strategy has been designed for multitask learning. TDP assigns each device with a single task and a single data loader while the total number of training samples remains fixed across all devices for each task.
(6) M3FM can be ”linearly” adapted to new tasks by modifying image linear embedding layer and task-specific predictors.
(7) M3FM serves as a medical research platform. This is evidenced by our extensive experimental results demonstrating that by synergizing multimodal data and multitasks, our M3FM model significantly improves the performance of the counterpart model trained on single-modality data for individual tasks, and the informative data types for each task can be identified through ablation inference or training. Also, M3FM can be adapted to achieve a decent performance of new tasks with a small out-of-distribution dataset.

### v3

@@ 4.2.4 Task Encoder
Figure 6-(e) illustrates the task encoder, designed to extract task-specific embedding features from the multimodal token embeddings, given the special $<TASK>$ token embedding. The task encoder was implemented with multiple Transformer layers and regards all tokens as a single input sequence. Note that only the special $<TASK>$ token is forwarded to the task encoder and the rest of the question tokens are ignored, as we empirically found that other tokens are not useful in practice. The special token embedding from the final Transformer layer serves as the task-specific embedding feature that integrates all multimodal data. For M3FM-Base, M3FM-Large, and M3FM-Huge, the number of Transformer layers is 4 in every case, and the sizes of task token embeddings are 768, 1,024, and 1,280, respectively.
@@ 4.3 Predictors
The Predictors map task-specific embedding features to answers. In this study, we found that the task-specific predictor can be automatically and perfectly selected through the Task ID Predictor, which takes the $<TASK>$ embedding corresponding to the question text. We implemented all Predictors including the Task ID Predictor as a two-layer MLP. Different tasks may have different Predictors or shared Predictors for the same output dimension, such as ”Yes” or ”No” answers. Similar to language models that regard text generation as the token classification over a vocabulary problem, we formulate our answer prediction as a classification problem over the predefined answer candidates, as summarized in Table 1, except for six-year lung cancer risk prediction. Following [33], we formulate lung cancer risk prediction as a hazard regression problem. It is worth mentioning that more types of prediction tasks, such as image segmentation and object detection, can be incorporated into M3FM by adding the corresponding lightweight task-specific predictors as demonstrated in our previous study [48].
@@ 4.4 Self-supervised Pretraining
A key step to optimize large models is self-supervised pretraining with large unlabeled datasets. In this study, we adapted the masked autoencoder method [40, 41] to pretrain our CTViT on the constructed OpenM3Chest pretraining dataset. Figure 6-(b) shows the CTViT pretraining architecture, which consists of CTViT, an image decoder, and a multi-scale linear de-tokenizer. The image encoder introduced in Subsection 4.2.2 was optimized by predicting masked cubes (85%) from a small number of visible cubes (15%). To reduce the memory overhead, only some selected slices along the longitude direction were predicted while recovering each 3D patch. We pre-trained CTViT with the defined multi-scale 3D CT volumes and a set of data augmentation operations, including random cropping, rotation, resizing, and perturbed display windowing. The text Transformer in M3FM was initialized with the off-the-shelf RoBERTa model pre-trained via masked language modeling and then trained end-to-end [49, 39].
@@ 4.5 Multitask Learning
After self-supervised pretraining, M3FM can be trained with any combination of different tasks with properly selected multimodal datasets by optimizing multitask loss functions simultaneously. We used the sigmoid cross-entropy loss function for the CVD mortality risk and lung cancer risk prediction tasks and the softmax cross-entropy loss function for all other tasks. As the number of tasks increases, there is a significant rise in computational cost. To address this problem, we design a distributed task-parallel (DTP) training strategy. TDP assigns each computing device with a single task and a single data loader while the total number of training samples remains fixed across all devices for each task. Since M3FM is a unified model capable of handling various tasks, despite differences in input and output dimensions, gradients computed across all tasks can be readily accumulated, enabling simultaneous parameter optimization.
@@ 4.6 Transfer Learning
M3FM is designed for adaptability and generalization, enabling the enhancement of out-of-distribution task performance through transfer learning. This capability extends to new tasks with varying image input dimensions, clinical data types, and output dimensions. To accommodate different image dimensions, the addition of a linear embedding layer suffices. For diverse clinical datasets, we can simply describe involved clinical data in free text to the model, without needing any modification on the M3FM architecture, as shown in Figure 5-(a). Specifically, adjusting to different output dimensions requires only the inclusion of a lightweight predictor. Consequently, M3FM can be easily fine-tuned to enable new tasks by leveraging the pre-trained model parameters.
4.7 Training Details
We used the AdamW optimizer [50], cosine decay learning rate schedule [51], weight decay of 0.05, and automatic mixed precision in PyTorch for training all models. In pre-training CTViT, the batch size was 192, the learning rate was $3.75\times 10^{-4}$, on each GPU a single scale of CT input was randomly selected from the four sizes described in Section 4.2.2, the CTViT was trained for 200K iterations with 10K warmup iterations, the decoder depth was 2, the voxel values were normalized within each cube in calculating the MSE loss function, and the CT inputs were randomly scaled by the factor of [0.5, 2], [0.6, 1.4], and [0.6, 1.4] in axial, coronal, and sagittal dimensions, and the physical sizes were recalculated accordingly. In training task-specific models including the transfer learning, the batch size was 12, the number of training iterations was 30K with 2K warmup iterations, the learning rate was $2\times 10^{-4}$, and the layer-wise learning rate decay of 0.95 was used. In multitask training, the total batch size was 972, including 12 samples for each of the 17 tasks and 768 samples for the clinical information retrieval tasks. All CT inputs had a random HU range perturbation, random rotation degrees, and random padding in training M3FM models, with the corresponding hyperparameters in training for different tasks described in Table S4 in the supplementary, and each clinical data element was randomly included with the probability of 0.8.
4.8 Hardware Requirement
All our models are trained on the AiMOS Supercomputer in the Center for Computational Innovation at Rensselaer Polytechnic Institute (https://docs.cci.rpi.edu/clusters/DCS_Supercomputer/). For CTViT pretraining and multi-task training, we used 192 NVIDIA Tesla V100 GPUs with 32 GiB of memory each, i.e., 6 GPUs per node $\times$ 32 nodes. For all single-task training and finetuning, we used 12 NVIDIA Tesla V100 GPUs with 32 GiB of memory each, i.e., 6 GPUs per node $\times$ 2 nodes.
[…]
Data Availability The raw data of NLST are available from the NIH National Cancer Institute (https://cdas.cancer.gov/nlst/). The raw data of MIDRC are available at https://www.midrc.org/. Our curated OpenM3Chest dataset will be publicly available upon the acceptance of the paper. Restrictions apply to the availability of anonymized patient data from WFUSM and MGH for this project with institutional permission and are, thus, not publicly available.
[…]
Code Availability Demo codes and data are available at https://github.com/niuchuangnn/M3FM, and the integrated database and optimized codes will be made open to the academic community upon the publication of our paper.
[…]
Supplementary information
The accompanying supplementary file contains four tables to show the predefined sentence format for each clinical data element (Table S1), the labeling strategy for each task (Table S2), the predefined questions for each task (Table S3), and the hyperparameters of data transformations for each task (Table S4).
[…]
Acknowledgments
The imaging and associated clinical data downloaded from MIDRC (The Medical Imaging and Data Resource Center) and used for research in this study was made possible by the National Institute of Biomedical Imaging and Bioengineering (NIBIB) of the National Institutes of Health under contract 75N92020D00021. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.